# An abstract framework for heterogeneous coupling: stability, approximation and applications

### v1

@@ 4. Solving the resulting discrete problem
Since we are interested in large problems, it is natural to consider an iterative solver for the reduced system of Problem 3.1, where the stiffness matrix relative to the approximate Schur complement operator need not be assembled, but where, at each iteration, the approximate black box solvers are called upon to evaluate their action on a given element of $\Lambda_{\delta}$. We note that for very large systems it may be advantageous to proceed in a nested fashion, where also the subproblems are themselves decomposed in smaller subproblems, until the local solvers are so small that direct solvers can be applied. Focusing on the unstabilized Problem 3.1, and following once again an approach inspired by the FETI method, in order to efficiently solve Problem 3.1 we will need
@@
    to reduce the saddle point problem posed on $\Lambda_{\delta}\times Z$ to a positive definite problem defined on the subspace $\widehat{\Lambda}_{\delta}$ of $\Lambda_{\delta}$, which can be solved by a Krylov type method (e.g. PCG, if the corresponding operator is symmetric)
@@
    to construct a preconditioner for the the restriction of the bilinear form $s_{h}$ to the subspace $\widehat{\Lambda}_{\delta}$.
The space $\widehat{\Lambda}_{\delta}$ and its dual
We will need, in the following, to simultaneously handle the dual spaces $\Lambda_{\delta}^{\prime}$ and $\widehat{\Lambda}_{\delta}^{\prime}$ of, respectively, $\Lambda_{\delta}$ and $\widehat{\Lambda}_{\delta}=\Lambda_{\delta}\cap\ker G^{T}$, and, in doing this, we need to take into account the relation between these two duals. To this aim, it will be useful to think of $\widehat{\Lambda}_{\delta}$ as an independent space and, to make this point of view explicit, we introduce the natural injection operator $\mathcal{I}:\widehat{\Lambda}_{\delta}\to\Lambda_{\delta}$. The space $\widehat{\Lambda}_{\delta}^{\prime}$ is isomorphic to the quotient space $\Lambda_{\delta}^{\prime}/\ker\mathcal{R}$ where $\mathcal{R}=\mathcal{I}^{T}:\Lambda_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}^{\prime}$ is the natural restriction operator, defined as
$\langle\mathcal{R}\phi,\widehat{\lambda}\rangle=\langle\phi,\mathcal{I}\widehat{\lambda}\rangle=\langle\phi,\widehat{\lambda}\rangle.$
The elements of $\widehat{\Lambda}_{\delta}^{\prime}$ are then naturally defined as equivalence classes ${\widetilde{\varphi}}=[\phi]$, $\phi\in\Lambda_{\delta}^{\prime}$, where $\phi^{\prime}\in[\phi]$ iff $\mathcal{R}(\phi-\phi^{\prime})=0$. In constructing our solver, we will need to identify elements of $\widehat{\Lambda}_{\delta}^{\prime}$ with a canonical representative $\phi\in\Lambda_{\delta}^{\prime}$ minimizing a suitable norm. To this aim we introduce a discrete scalar product $\sigma:\Lambda_{\delta}\times\Lambda_{\delta}\to\mathbb{R}$, not necessarily equivalent to the $\Lambda$ scalar product, and we define a mapping $\mathcal{K}:\widehat{\Lambda}_{\delta}^{\prime}\to\Lambda_{\delta}^{\prime}$ with, for all ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$, ${\widetilde{\varphi}}=[\mathcal{K}{\widetilde{\varphi}}]$, defined as
@@
$\mathcal{K}{\widetilde{\varphi}}=\mathop{\text{argmin}}_{{\phi^{\prime}\in\Lambda_{\delta}^{\prime}}\ :\ {{\widetilde{\varphi}}=[\phi^{\prime}]}}\|\phi^{\prime}\|_{\sigma,*},\qquad\text{ where }\quad\|\phi^{\prime}\|_{\sigma,*}=\sup_{\lambda\in\Lambda_{\delta}}\frac{\langle\phi^{\prime},\lambda\rangle}{\|\lambda\|_{\sigma}}.$
We have the following proposition clarifying the relation between the canonical representative $\mathcal{K}({\widetilde{\varphi}})$ of an element ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$ and any element $\phi\in\Lambda_{\delta}^{\prime}$ such that $[\phi]={\widetilde{\varphi}}$.
Proposition 4.1.
Let $\Pi_{\sigma}:\Lambda_{\delta}\to\widehat{\Lambda}_{\delta}$ denote the $\sigma$-orthogonal projector onto $\widehat{\Lambda}_{\delta}$. Then for all $\phi\in\Lambda_{\delta}^{\prime}$ it holds
$\mathcal{K}([\phi])=\Pi_{\sigma}^{T}\phi.$

### v2

@@ 4. Solving the resulting discrete problem
Since we are interested in large problems, it is natural to consider an iterative solver for the reduced system of Problem 3.1, where the stiffness matrix relative to the approximate Schur complement operator needs not be assembled, but where, at each iteration, the approximate black box solvers are called upon to evaluate their action on a given element of $\Lambda_{\delta}$. We note that for very large systems it may be advantageous to proceed in a nested fashion, where also the subproblems are themselves decomposed in smaller subproblems, until the local solvers are so small that direct solvers can be applied. Focusing on the unstabilized Problem 3.1, and following once again an approach inspired by the FETI method, in order to efficiently solve Problem 3.1 we will need to
@@
    reduce the saddle point problem posed on $\Lambda_{\delta}\times Z$ to a positive definite problem, defined on the subspace $\widehat{\Lambda}_{\delta}$ of $\Lambda_{\delta}$, which can be solved by a Krylov type method (e.g. PCG, if the corresponding operator is symmetric);
@@
    construct a preconditioner for the the restriction of the bilinear form $s_{h}$ to the subspace $\widehat{\Lambda}_{\delta}=\Lambda_{\delta}\cap\widehat{\Lambda}$.
To this aim, we will leverage a number of linear operators that we will have to construct either directly or through the corresponding bilinear form.
Reduction to $\widehat{\Lambda}_{\delta}$
In order to reduce the problem to the subspace $\widehat{\Lambda}_{\delta}$, we start by constructing a mapping $\mathcal{K}:\widehat{\Lambda}_{\delta}^{\prime}\to\Lambda_{\delta}^{\prime}$ which will also be instrumental in the definition of the preconditioner. To this aim, letting $G:Z\to\Lambda_{\delta}^{\prime}$ be defined, with a slight abuse of notation, as
(4.1) $\langle Gz,\lambda\rangle=\langle B_{\delta}z,\lambda\rangle=b(z,\lambda),\ \forall z\in Z,\ \lambda\in\Lambda_{\delta},$
we start by choosing an easily computable discrete symmetric positive semi definite bilinear form $\sigma^{*}:\Lambda_{\delta}^{\prime}\times\Lambda_{\delta}^{\prime}\to\mathbb{R}$, which we assume to be continuous on $\Lambda_{\delta}^{\prime}$ and coercive on $\mathfrak{R}(G)$: for all $z\in Z$, $\phi,\psi\in\Lambda_{\delta}^{\prime}$
(4.2) $\sigma^{*}(Gz,Gz)\gtrsim\|Gz\|_{\Lambda^{\prime}}^{2},\qquad\sigma^{*}(\phi,\psi)\lesssim\|\phi\|_{\Lambda^{\prime}}\|\psi\|_{\Lambda^{\prime}}.$
The bilinear form $\sigma^{*}$ induces a scalar product on $\mathfrak{R}(G)$, which allows to define a projector $Q_{\sigma}^{*}:\Lambda_{\delta}^{\prime}\to\mathfrak{R}(G)$ (easily computable) as
(4.3) $Q_{\sigma}^{*}\phi\in\mathfrak{R}(G)\qquad\text{ solution of }\qquad\sigma^{*}(Q_{\sigma}^{*}\phi,\psi)=\sigma^{*}(\phi,\psi)\qquad\forall\psi\in\mathfrak{R}(G).$
Remark 4.1.
Choosing $\sigma^{*}$ equal to the $\Lambda^{\prime}$ scalar product automatically yields (4.2). In principle, we could then use such a scalar product in the place of the bilinear form $\sigma^{*}$. $Q_{\sigma}^{*}$ would then be the $\Lambda^{\prime}$ orthogonal projection onto $\mathfrak{R}(G)$. However, in our framework, such a choice would often results in $Q_{\sigma}^{*}$ being quite difficult to compute. This is the reason why a simpler bilinear form $\sigma^{*}$ needs to be introduced.
Let now ${\widetilde{\varphi}}\in\widehat{\Lambda}_{\delta}^{\prime}$. We define $\mathcal{K}{\widetilde{\varphi}}\in\Lambda_{\delta}^{\prime}$ by setting
(4.4) $\mathcal{K}{\widetilde{\varphi}}=\phi-Q_{\sigma}^{*}\phi,$
where $\phi\in\Lambda_{\delta}^{\prime}$ is any element such that
@@
$\langle{\widetilde{\varphi}},\widehat{\lambda}\rangle=\langle\phi,\widehat{\lambda}\rangle,\qquad\text{ for all }\widehat{\lambda}\in\widehat{\Lambda}_{\delta}.$
It is not difficult to check that if $\langle\phi-\phi^{\prime},\widehat{\lambda}\rangle=0$ for all $\widehat{\lambda}\in\widehat{\Lambda}_{\delta}$, then $\phi-Q_{\sigma}^{*}\phi=\phi^{\prime}-Q_{\sigma}^{*}\phi^{\prime}$, so that $\mathcal{K}$ is well defined, independently on how $\phi$ is chosen.
Remark 4.2.
If we think of $\widehat{\Lambda}_{\delta}$ and $\Lambda_{\delta}$ as independent spaces, and introduce an operator $\mathcal{I}:\widehat{\Lambda}_{\delta}\to\Lambda_{\delta}$, representing the natural injection, we can identify $\widehat{\Lambda}_{\delta}^{\prime}$ with the quotient space $\Lambda_{\delta}^{\prime}/\ker\mathcal{R}$ where $\mathcal{R}=\mathcal{I}^{T}:\Lambda_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}^{\prime}$ is the natural restriction operator, defined as
$\langle\mathcal{R}\phi,\widehat{\lambda}\rangle=\langle\phi,\mathcal{I}\widehat{\lambda}\rangle\qquad\phi\in\Lambda_{\delta}^{\prime},\ \widehat{\lambda}\in\widehat{\Lambda}_{\delta}.$
It is possible to show that the equivalence class $\phi]$ of an element $\phi\in\Lambda_{\delta}^{\prime}$ can be identified with $\phi+\mathfrak{R}(G)$. The operator $\mathcal{K}:\widehat{\Lambda}_{\delta}^{\prime}\to\Lambda_{\delta}^{\prime}$ can be interpreted as the selection of a specific “canonical” representative in $\Lambda_{\delta}^{\prime}$ of an equivalence class ${\widetilde{\varphi}}=[\phi]\simeq\phi+\mathfrak{R}(G)$ in $\widehat{\Lambda}_{\delta}^{\prime}$. Remark that $\mathcal{K}$ can be naturally extended to a mapping $\mathcal{K}:\Lambda_{\delta}^{\prime}\to\Lambda_{\delta}^{\prime}$ that, given any element $\phi$ of $\Lambda_{\delta}^{\prime}$ selects the canonical representative of the equivalence class $\phi]=\phi+\mathfrak{R}(G)$.
Remark 4.3.
One possible way of constructing a bilinear form $\sigma^{*}$ with the desired properties is by duality with respect to a scalar product on the space $\Lambda_{0}$ given by Assumption 3.8. Indeed, let $\sigma:\Lambda_{0}\times\Lambda_{0}\to\mathbb{R}$ be a scalar product inducing on $\Lambda_{0}$ a norm equivalent to $\|\cdot\|_{\Lambda}$. More precisely, assume that for all $\lambda,\mu\in\Lambda_{0}$, $\sigma$ satisfies
(4.5) $\sigma(\lambda,\lambda)\gtrsim\|\lambda\|^{2}_{\Lambda},\qquad\sigma(\lambda,\mu)\lesssim\|\lambda\|_{\Lambda}\|\mu\|_{\Lambda}.$
Letting $\{\eta_{i},\ i=1,\cdots,N\}$ be a basis for $\Lambda_{0}$, $\Sigma=(\sigma_{ij})_{ij}$, with $\sigma_{ij}=\sigma(\eta_{j},\eta_{i})$, the stiffness matrix induced by $\sigma$, and $\Sigma^{-1}=(\sigma^{*}_{ij})_{ij}$ its inverse, we can define $\sigma^{*}$ as
$\sigma^{*}(\phi,\psi)=\sum_{i,j=1}^{N}\sigma^{*}_{ij}\langle\phi,\eta_{j}\rangle\langle\psi,\eta_{i}\rangle.$
By the arguments in [9] it is not difficult to check that $\sigma^{*}$ is a scalar product on $\mathfrak{R}(G)$ inducing a norm equivalent to the $\Lambda^{\prime}$ norm, and that the corresponding projector $Q^{*}_{\sigma}$ satisfies
$\|Q^{*}_{\sigma}\phi\|_{\Lambda^{\prime}}\lesssim\|\phi\|_{\Lambda^{\prime}},$
the implicit constant in the inequality only depending on the coercivity and continuity constants implicit in (4.5).
We let $\Pi_{\sigma}:\Lambda_{\delta}\to\Lambda_{\delta}$ be defined as
(4.6) $\Pi_{\sigma}=(\mathbf{1}_{{\Lambda_{\delta}^{\prime}}}-Q_{\sigma}^{*})^{T}=\mathbf{1}_{{\Lambda_{\delta}}}-(Q_{\sigma}^{*})^{T}.$
It is not difficult to check that $\mathfrak{R}(\Pi_{\sigma})\subseteq\widehat{\Lambda}_{\delta}$ and that $\lambda_{\delta}\in\widehat{\Lambda}_{\delta}$ implies that $\Pi_{\sigma}\lambda_{\delta}=\lambda_{\delta}$. Then $\Pi_{\sigma}$ is a projection onto $\widehat{\Lambda}_{\delta}$. Letting $\lambda$ be the solution to Problem 3.1, we can see that $\lambda^{0}=\lambda-\Pi_{\sigma}\lambda$ can be computed as
$\lambda^{0}=(Q_{\sigma}^{*})^{T}Gz^{0},\quad\text{ with $z^{0}\in Z$ solution of }\quad\sigma^{*}(Gz^{0},Gz)=-\langle f,z\rangle,\ \forall z\in Z,$
and that $\widehat{\lambda}=\Pi_{\sigma}\lambda=\lambda-\lambda^{0}$ is the solution to the following reduced problem
We know by Lemma 3.6 that $s_{h}$ is coercive on $\widehat{\Lambda}_{\delta}$, and, consequently, that the corresponding operator $\widehat{S}_{h}:\widehat{\Lambda}_{\delta}\to\widehat{\Lambda}_{\delta}^{\prime}$ is invertible. The reduced problem (4.7) can be then addressed by any iterative solver well suited to solve linear systems with positive definite matrices. For instance, if, in addition, the operator $A^{+}_{h}$ is symmetric, which implies that $S_{h}$ and, consequently, $\widehat{S}_{h}$ are symmetric, one can resort to the preconditioned conjugate gradient method.
4.1. Construction of the preconditioner $\widehat{M}^{-1}$
A preconditioner $\widehat{M}^{-1}:\widehat{\Lambda}_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}$ will be constructed by combining the mapping $\mathcal{K}$ with a suitably defined preconditioning operator $M^{+}:\Lambda_{\delta}^{\prime}\to\Lambda_{\delta}$, corresponding to a bilinear form $m^{+}:\Lambda_{\delta}^{\prime}\times\Lambda_{\delta}^{\prime}\to\mathbb{R}$. We will build $\widehat{M}^{-1}$ by building the corresponding bilinear form $m^{*}:\widehat{\Lambda}_{\delta}^{\prime}\times\widehat{\Lambda}_{\delta}^{\prime}\to\mathbb{R}$ such that
$\langle\widehat{M}^{-1}{\widetilde{\varphi}},\widetilde{\psi}\rangle=m^{*}({\widetilde{\varphi}},\widetilde{\psi}),\qquad{\widetilde{\varphi}},\widetilde{\psi}\in\widehat{\Lambda}_{\delta}^{\prime}.$
The bilinear form $m^{*}$ will have the form
[…]
where $m^{+}:\Lambda_{\delta}^{\prime}\times\Lambda_{\delta}^{\prime}\to\mathbb{R}$, which is the bilinear form that we we will actually construct, corresponds to a linear operator $M^{+}:\Lambda_{\delta}^{\prime}\to\Lambda_{\delta}$ “mimicking” some kind of inverse of $S_{h}$, if $S_{h}$. If $A^{+}_{h}$ and $B_{\delta}$ were invertible, this would be
(4.8) $(BA^{+}_{h}B^{T})^{-1}=B^{-T}(A^{+}_{h})^{-1}B^{-1}.$
Of course, none of the above mentioned operators is generally invertible; the idea is then to use the right-hand side expression in (4.8) as a guideline, and replace $B^{-1}$ and $(A_{h}^{+})^{-1}$, respectively, with a suitable pseudo inverse $B_{\delta}^{+}$, and with $A$ (or some spectrally equivalent $\widetilde{A}$).
Following the approach put forward by the FETI preconditioner, we rely on three ingredients. The first ingredient is a finite dimensional space $V_{\delta}\subseteq V$, to which the image of $B_{\delta}^{+}$ will belong. We assume $V_{\delta}$ to satisfy
(4.9) $Z\subseteq V_{\delta},\qquad\text{and}\qquad\inf_{\lambda\in\Lambda_{\delta}}\sup_{v\in V_{\delta}}\frac{b(v,\lambda)}{\|v\|_{V}\|\lambda\|_{\Lambda}}\gtrsim 1.$
Observe that, if a space $V_{h}$ underlying the approximate pseudo inverse $A^{+}_{h}$ is available, a natural choice for $V_{\delta}$ is $V_{\delta}=V_{h}$, but, in our framework, we are also interested in situations in which such a space might not be accessible. Also note that $V_{\delta}$ needs not satisfy any approximation assumption, so that, depending on the framework, its dimension can be sensibly smaller that the dimension of $V_{h}$. We let $A_{\delta}:V_{\delta}\to V_{\delta}^{\prime}$ and $B_{\delta}:V_{\delta}\to\Lambda_{\delta}^{\prime}$, be the discrete versions of the operators $A$ and $B$, defined as
$\langle A_{\delta}v,w\rangle=a(v,w)\ \forall v,w\in V_{\delta},\qquad\langle B_{\delta}v,\lambda\rangle=b(v,\lambda)\ \forall v\in V_{\delta},\lambda\in\Lambda_{\delta}.$
The second ingredient is a discrete positive semi definite bilinear form $d:V_{\delta}\times V_{\delta}\to\mathbb{R}$ such that $d$ is coercive on $\ker B_{\delta}$, and a corresponding splitting of the space $V_{\delta}$ as
$V_{\delta}=\ker B_{\delta}\oplus(\ker B_{\delta})^{\perp},$
where
$\ker B_{\delta}=\{v\in V_{\delta}:\ b(v,\mu)=0,\ \forall\mu\in\Lambda_{\delta}\},$
$(\ker B_{\delta})^{\perp}=\{v\in V_{\delta}:\ d(v,w)=0,\ \forall w\in\ker B_{\delta}\}.$
We let $Q_{d}:V_{\delta}\to\ker B_{\delta}$ denote the $d$-orthogonal projection onto $\ker B_{\delta}$, and $\Pi_{d}:V_{\delta}\to(\ker B_{\delta})^{\perp}$ be defined as $\Pi_{d}=\mathbf{1}_{{V_{\delta}}}-Q_{d}$. We observe that for all $w\in V_{\delta}$ we have that
(4.10) $d(Q_{d}w,Q_{d}w)\leq d(w,w),\qquad d(\Pi_{d}w,\Pi_{d}w)\leq d(w,w).$
Whichever the choice of the bilinear form $d$, the restriction of $B_{\delta}$ to $(\ker B_{\delta})^{\perp}$ is injective and, thanks to the inf-sup condition (4.9), surjective. We then let $B_{\delta}^{+}:\Lambda_{\delta}^{\prime}\to(\ker B_{\delta})^{\perp}$ denote its inverse, which is a pseudo inverse of $B_{\delta}$, satisfying
We easily realize that, given $\phi\in\Lambda_{\delta}^{\prime}$, the pseudo inverse $B_{\delta}^{+}\phi$ can be defined as $B_{\delta}^{+}\phi=v$, where $(v,\zeta)\in V_{\delta}\times\Lambda_{\delta}$ is the uniquely defined solution to
(4.12) $\begin{array}[]{lcl}d(v,w)+b(w,\zeta)&=&0,\qquad\forall w\in V_{\delta}\\ b(v,\mu)&=&\langle\phi,\mu\rangle,\qquad\forall\mu\in\Lambda_{\delta}.\end{array}$
The third ingredient is a set of local preconditioners for $A_{\delta}$: for all $k$ we let $\widetilde{a}^{k}:V_{\delta}^{k}\times V_{\delta}^{k}\to\mathbb{R}$ denote a coercive bilinear form satisfying, for some constants $K^{*}\geq 1\geq\kappa_{*}>0$,
(4.13) $\kappa_{*}a^{k}(v,v)\leq\widetilde{a}^{k}(v,v)\leq K^{*}a^{k}(v,v)\qquad\forall v\in V_{\delta}^{k}.$
We let $\widetilde{A}_{\delta}^{k}$ denote the corresponding linear operator, and we let $\widetilde{a}:V_{\delta}\times V_{\delta}\to\mathbb{R}$ and $\widetilde{A}_{\delta}:V_{\delta}\to V_{\delta}^{\prime}$ denote the related product bilinear form and corresponding linear operator. We let $|\cdot|_{\widetilde{a}}$ denote the seminorm induced on $V_{\delta}$ by the bilinear form $\widetilde{a}$:
(4.14) $|v|_{\widetilde{a}}=\sqrt{\widetilde{a}(v,v)}.$
Following the guidelines provided by the formal inversion expression (4.8) we introduce the linear operator $M^{+}:\Lambda_{\delta}^{\prime}\to\Lambda_{\delta}$, defined as
$M^{+}=(B_{\delta}^{+})^{T}\widetilde{A}_{\delta}B_{\delta}^{+},$
which corresponds to the bilinear form $m^{+}:\Lambda_{\delta}^{\prime}\times\Lambda_{\delta}^{\prime}$ defined as
$m^{+}(\phi,\psi)=\widetilde{a}(B_{\delta}^{+}\phi,B_{\delta}^{+}\psi).$
By combining the bilinear form $m^{+}$ with the mapping $\mathcal{K}$ defined by (4.4), we finally build our preconditioner $\widehat{M}^{-1}:\widehat{\Lambda}_{\delta}^{\prime}\to\widehat{\Lambda}_{\delta}$ as the operator corresponding to the bilinear form $m^{*}:\widehat{\Lambda}_{\delta}^{\prime}\times\widehat{\Lambda}_{\delta}^{\prime}\to\mathbb{R}$ defined as
[…]
We have the following theorem, the proof of which is fundamentally the same as the proof of the analogous result for the FETI method in the domain decomposition framework (see [42]). For the sake of completeness we give the proof in an appendix.
Theorem 4.4.
Let $\Pi_{d}:V_{\delta}\to V_{\delta}$ denote the d{d}-orthogonal projection operator onto the d{d}-orthogonal complement $(\ker B_{\delta})^{\perp}$ of $\ker B_{\delta}$ in $V_{\delta}$ and, for $v\in V_{\delta}$ and $L:V_{\delta}\to V_{\delta}$ linear operator, let
$\|v\|_{\widetilde{a}}^{2}=|v|_{\widetilde{a}}^{2}+\kappa_{*}\|\bar{v}\|^{2}_{V},\qquad\|L\|_{\widetilde{a}\to\widetilde{a}}=\sup_{w\in V_{\delta}}\frac{\|Lw\|_{\widetilde{a}}}{\|w\|_{\widetilde{a}}}$,
$\bar{v}\in Z$ denoting the $V$-orthogonal projection of $v$. Then it holds
$\kappa(\widehat{M}^{-1}\widehat{S}_{h})\lesssim C(\widetilde{a},V_{\delta},d,\sigma)$}
with
$C(\widetilde{a},V_{\delta},d,\sigma)=\frac{K^{*}}{\kappa_{*}}\|\Pi_{d}\|_{\widetilde{a}\to\widetilde{a}}^{2}\left(1+\|B_{\delta}^{+}Q_{\sigma}^{*}B_{\delta}\|_{\widetilde{a}\to\widetilde{a}}\right)^{2}.$
Remark 4.5.
As, thanks to (4.13), $|v|_{\widetilde{a}}=0$ if and only if $v\in Z$, $\|\cdot\|_{\widetilde{a}}$ is indeed a norm on $V_{\delta}$.
4.2. Efficiency of the preconditioner
While the operators $A_{\delta}$, $B_{\delta}$ and $G$ are substantially given by the continuous problem, there is, a priori, a certain freedom in the choice of the space $V_{\delta}$ and of the of the bilinear forms $\sigma^{*}$ and $d$. These have to be chosen in such a way that the resulting preconditioner $\widehat{M}^{-1}$ is cheap (or easily parallelizable), and that, at the same time, the norms of $\Pi_{d}$ and $B_{\delta}^{+}(Q_{\sigma}^{*})^{T}B_{\delta}$ stay bounded. How this can be done, and whether this can be done efficiently, depends on the properties of the actual problem considered. Giving a general abstract recipe simultaneously allowing for efficient parallel implementation and optimal or quasi optimal condition number bounds, by solely relying on the assumptions made up to now is, in our opinion, not feasible. We can, nevertheless, guarantee efficiency of our abstract preconditioning strategy, provided the continuous multiplier space $\Lambda$ has some suitable additional localization property. More precisely, we make the following assumption.
Assumption 4.6.
There exist linearly independent subspaces $\Lambda^{\ell}\subseteq\Lambda$ such that
  1. (1)
    $\oplus\Lambda^{\ell}$ is dense in $\Lambda$;
  2. (2)
    for all $\ell$ there exist $k^{+}_{\ell}$}, $k^{-}_{\ell}$} such for all $k\not\in\{{k^{+}_{\ell}},{k^{-}_{\ell}}\}$, $v\in V_{k}$ and $\lambda\in\Lambda^{\ell}$ imply $b_{k}(v,\lambda)=0$.
Essentially, Assumption 4.6 requires that, at the continuous level, (a dense subspace of) the multiplier space can be decomposed in subspaces each “seeing” at most two subproblems. Under such an assumption it is always possible to choose $\Lambda_{\delta}$ as $\Lambda_{\delta}=\oplus\Lambda_{\delta}^{\ell}$, $\Lambda_{\delta}^{\ell}\subseteq\Lambda^{\ell}$. We remark that such an assumption could be relaxed by allowing each $\Lambda^{\ell}$ to “see” at most $N$ subspaces $V^{k}$, but for the sake of notational simplicity we do not address such an option, which can however be tackled by a similar strategy.
We let $\partial(k)=\{\ell:\ {k^{+}_{\ell}}=k\ \text{ or },{k^{-}_{\ell}}=k\}$ denote the indexes of all the subspaces $\Lambda^{\ell}$ that “see” the space $V_{\delta}^{k}$, and $\mathcal{N}(\ell)=\partial({k^{+}_{\ell}})\cup\partial({k^{-}_{\ell}})$ denote the indexes of all the $\Lambda^{\ell^{\prime}}$ that share with $\Lambda^{\ell}$ a neighboring space $V_{h}^{k}$. For all $k$, we introduce subspaces $V^{k}_{0}=\ker B_{\delta}^{k}$ and
$V^{k}_{\ell}=\{v\in V_{\delta}^{k},\ \widetilde{a}^{k}(v,w)=0,\ \forall w\in V^{k}_{0}\ \text{ and }\ b^{k}(v,\mu)=0,\ \forall\mu\in\Lambda^{i},\ \forall i\not$$\ell\}$.
We make the following assumptions on the space $V_{\delta}$
Assumption 4.7.
For all $k$ the space $V_{\delta}^{k}$ can be split as
$V_{\delta}^{k}=V^{k}_{0}\oplus_{\ell\in\partial(k)}V^{k}_{\ell}$
and there exist constants $c^{\flat}$ and $C^{\flat}$ such that for all $k$, for all $v^{k}=v^{k}_{0}\oplus_{\ell\in\partial(k)}v^{k}_{\ell}$ it holds
$c^{\flat}\|v^{k}\|_{\widetilde{a}}^{2}\leq\|v^{k}_{0}\|_{\widetilde{a}}^{2}+\sum_{\ell\in\partial(k)}\|v^{k}_{\ell}\|_{\widetilde{a}}^{2}\leq C^{\flat}\|v^{k}\|_{\widetilde{a}}^{2}.$
We remark that, by triangle inequality, $c^{\flat}$ is bounded from below by a constant depending only on the cardinality of $\partial(k)$, while the constant $C^{\flat}$ depends on the properties of the norms, bilinear forms and spaces involved. In particular, in the framework of the FETI method for domain decomposition, $C^{\flat}$ behaves like $\log(\delta/H)^{2}$, $\delta$ and $H$ denoting respectively the meshsize and the diameter of the subdomains. Under these assumptions we can define the scalar product $d$ as
(4.15) $d(u,v)=\sum_{k}d^{k}(u^{k},v^{k}),\qquad d^{k}(u^{k},v^{k})=\widetilde{a}^{k}(u^{k}_{0},v^{k}_{0})+\sum_{\ell\in\partial(k)}\widetilde{a}^{k}(u^{k}_{\ell},v^{k}_{\ell}).$
We claim that, for such a definition of $d$, the pseudo inverse $B_{\delta}^{+}$ can be evaluated by solving, in parallel, a set of “local” problems. Indeed, we have the following lemma.
Lemma 4.8.
Let $\phi\in\Lambda_{\delta}^{\prime}$ be given, and for all $\ell$ let $(v_{\ell}^{+},v_{\ell}^{-})\in V^{k^{+}_{\ell}}_{\ell}\times V^{k^{-}_{\ell}}_{\ell}$, $\zeta_{\ell}\in\Lambda^{\ell}$ be the solution to
(4.16) $\widetilde{a}^{k^{+}_{\ell}}(v_{\ell}^{+},w^{+})-b^{k^{+}_{\ell}}(w^{+},\zeta_{\ell})=0,\qquad\forall w^{+}\in V^{k^{+}_{\ell}}_{\ell},$
(4.17) $\widetilde{a}^{k^{-}_{\ell}}(v_{\ell}^{-},w^{-})-b^{k^{-}_{\ell}}(w^{-},\zeta_{\ell})=0,\qquad\forall v^{-}\in V^{k^{-}_{\ell}}_{\ell},$
(4.18) $b^{k^{+}_{\ell}}(v_{\ell}^{+},\mu)+b^{k^{-}_{\ell}}(v^{-}_{\ell},\mu)=\langle\phi,\mu\rangle,\qquad\forall\mu\in\Lambda^{\ell}.$
Then, letting $v_{\ell}\in V_{h}$ be defined by
$v_{\ell}^{k^{+}_{\ell}}=v_{\ell}^{+},\qquad v_{\ell}^{k^{-}_{\ell}}=v_{\ell}^{-},\qquad v_{\ell}^{i}=0,\quad i\not\in\{{k^{+}_{\ell}},{k^{-}_{\ell}}\}$
it holds that
(4.19) $B^{+}\phi=\sum_{\ell}v^{\ell}.$